OpenStudy (anonymous):

Barack Obama and Steve Jobs had a race. Obama running speed was 7 m/s while Job's was 8.5 m/s. They started off at the same time. How far apart were they after 8 seconds?

OpenStudy (bahrom7893):

Obama won the race, that's all I know

OpenStudy (bahrom7893):

the question is wrong, obama's supposed to be faster

OpenStudy (bahrom7893):

anyway all jokes aside..

OpenStudy (anonymous):

your bad.

OpenStudy (bahrom7893):

in 8 seconds, obama ran: 7*8 = 56m in 8 seconds, jobs ran: 8.5*8 = 68m

OpenStudy (bahrom7893):

68-56 = 12, so they were 12 m apart